OpenStudy (anonymous):

Solve: 3/2x + 1/5 = 3/10

jimthompson5910 (jim_thompson5910):

The equation is \[\large \frac{3}{2}x + \frac{1}{5} = \frac{3}{10}\] right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

multiply both sides by the LCD 10 to get \[\large \frac{3}{2}x + \frac{1}{5} = \frac{3}{10}\] \[\large 10\left(\frac{3}{2}x + \frac{1}{5}\right) = 10\left(\frac{3}{10}\right)\] \[\large 10\left(\frac{3}{2}x\right) + 10\left(\frac{1}{5}\right) = 10\left(\frac{3}{10}\right)\] \[\large \frac{10}{1}\left(\frac{3}{2}x\right) + \frac{10}{1}\left(\frac{1}{5}\right) = \frac{10}{1}\left(\frac{3}{10}\right)\] \[\large \frac{10*3}{1*2}x + \frac{10*1}{1*5} = \frac{10*3}{1*10}\] \[\large \frac{30}{2}x + \frac{10}{5} = \frac{30}{10}\] \[\large 15x + 2 = 3\] I'll let you take over

OpenStudy (anonymous):

So x= 1/15 ?

jimthompson5910 (jim_thompson5910):

yes that is correct

OpenStudy (anonymous):

Awsome! Thank you!

jimthompson5910 (jim_thompson5910):

np